# Contrastive Pretraining for Visual Concept Explanations of Socioeconomic Outcomes

### v1

@@ Abstract
Predicting socioeconomic indicators from satellite imagery with deep learning has become an increasingly popular research direction. Post-hoc concept-based explanations can be an important step towards broader adoption of these models in policy-making as they enable the interpretation of socioeconomic outcomes based on visual concepts that are intuitive to humans. In this paper, we study the interplay between representation learning using an additional task-specific contrastive loss and post-hoc concept explainability for socioeconomic studies. Our results on two different geographical locations and tasks indicate that the task-specific pretraining imposes a continuous ordering of the latent space embeddings according to the socioeconomic outcomes. This improves the model’s interpretability as it enables the latent space of the model to associate urban concepts with continuous intervals of socioeconomic outcomes. Further, we illustrate how analyzing the model’s conceptual sensitivity for the intervals of socioeconomic outcomes can shed light on new insights for urban studies.
@@ 1 Introduction
In this work, we introduce a new expandability pipeline (visualized in Figure 1) that uses post-hoc concept explanations to reveal how deep learning models predict socioeconomic outcomes in terms of easy-to-understand urban concepts. It consists of three main steps: 1) a task-specific contrastive pre-training of a deep learning encoder to continuously order the latent space embeddings based on their socioeconomic outcome, 2) probing a linear layer on the top of the encoder to predict socioeconomic outcome, and 3) concept testing with the Testing with Concept Activation Vectors (TCAV) method [17]. We evaluate our pipeline on two tasks for estimating socioeconomic indicators from remote sensing imagery, namely income prediction in France [1], and liveability estimation in the Netherlands [21]. In summary, our work evaluates for the first time visual post-hoc concept-based explanations in remote sensing socioeconomic studies with the following main contributions:
@@ 2 Related Work
Although essential for their wider adoption in practice, the interpretability of the deep learning models for socioeconomic indicator estimation is rarely tackled [14]. The approaches in related studies mainly propose using saliency maps [1] or concept bottleneck models [25, 20, 21]. Concretely, Abitbol and Karsai trains an EfficientNet deep learning model on aerial imagery to predict household income in five French cities [1]. They used the Grad-CAM saliency map method to infer the relevant regions for the model predictions. The identified regions are matched with the urban class polygons to reveal the relationship between urban classes and the income prediction of the deep learning model. Their results show that residential areas can both activate in case of high or low socioeconomic status depending on the city and area, whereas the presence of infrastructure such as motor- or railways is consistently related to a lower household income. When it comes to concept bottleneck models, they are used by Scepanovic et al. to predict the urban vitality in six Italian cities from Sentinel-2 imagery [25]. In the first step, the authors use a VGG-16 network and a convolutional autoencoder to predict six different land use and building block characteristics from image patches. These features are then used as input to the linear model that estimates urban vitality. The authors discover, that intersection density, small parks, anisotropicity, and land use mix positively influence vitality, while the variables block size and building height negatively impact the vitality. Further, Levering et al. train an end-to-end convolutional neural network with a semantic bottleneck to predict the landscape scenicness in the UK based on Sentinel-2 imagery [20]. The interpretable concepts in the semantic bottleneck layer encode the region’s land cover distribution. This architecture enables an explanation of the predicted scenicness score in terms of the land cover characteristics of the region. The authors find, that scenic areas are often associated with mountainous, coastal regions, while urban areas and the presence of human influence are negatively related to the landscape scenicness. The semantic bottleneck models are also used in [21] to predict the livability in the Dutch cities using aerial imagery. Again, the authors include a semantic bottleneck layer, which predicts five different intermediate domain scores. Their results indicate, that some intermediate domains such as buildings or physical environments are much better suited to be predicted from aerial imagery than others like safety or amenities.
@@ 4.3 Inference and Model Training
We predict the socioeconomic outcomes using a Resnet-50 encoder [15] for feature extraction whose output is fed to a linear layer that regresses the socioeconomic outcome. To train the model, we followed a similar training procedure as in [32]: First, we optimize the encoder for 400 iterations to minimize the RNC loss (the hyperparameter details are provided in Section 7, Appendix). Next, we freeze the weights of the feature encoder and train a single linear layer that uses the encoder output to predict the socioeconomic outcomes. The weights of the linear layer were optimized for 100 epochs with the $\mathcal{L}_{1}$ loss and we selected the model yielding the best $R^{2}$ score on the validation set. Our code is provided at github.com/IvicaObadic/rnc-4-visual-concept-explanations.
@@ 5.1 Prediction of Socioeconomic Indicators
To better understand the effect of task-specific pretraining on the latent space, in Figure 2 we use the t-distributed Stochastic Neighbor Embedding (t-SNE) method [28] to visualize the instance activations after the average pooling layer of the ResNet-50 encoder. The plots in the first column illustrate that the contrastive pre-trained encoder increasingly orders the instances in the latent space in an approximately 1D manifold according to their target label for both datasets. In contrast, the plots visualized in the second column depict that the models optimized to directly predict the regression score seem to result in a more scattered embedding space that fails to capture the continuous nature of the data, as there are multiple clusters of instances sharing similar target values located in various subspaces of the latent space. Finally, we note that for the task-specific pre-trained encoder, the liveability values typically display higher variance in their score compared to the income values for a local area in the embedding space. This can be attributed to the difference in complexity between these two tasks. The liveability reference data is based on 100 factors spanning 5 domains, from which liveability scores are derived [21]. As it is a composite score, the ranking difference between pairs will depend on more complex patterns, which might explain the comparable results of the contrastive pretraining with the supervised trained encoder on the liveability task.
Figure 2: Household income instance activations in the average pooling layer visualized with t-SNE. By pre-training with Rank-n Contrast, the latent space can be ordered according to the task regression values, rather than the visual features of images. This results in an embedding space that aligns with the socioeconomic outcomes, and therefore better suited for interpretability.
@@ 5.2 Concept Explanations
As described in Section 3.2, an essential step in TCAV is learning a concept activation vector (CAV) that encodes the direction in the latent space where the examples of a concept are located. In our study, we learn the CAVs per layer in the latent space of the ResNet-50 encoder by training a one-vs-all linear SVM classifier to discriminate the examples of the concept from the examples of all other concepts based on their layer activations. Figure 3 shows that while the concepts are already mostly linearly separable in the average pooling layer of the baseline encoder trained with the $L_{1}$ loss, enforcing a continuously ordered latent space with the Rank-N-Contrast pre-trained encoder enables close-to-perfect discrimination, with substantial improvements in the medium residential concept for the liveability dataset and impervious surfaces, vegetation, agriculture, and sparse residential concepts in both datasets.
Refer to captionRefer to caption
@@
Figure 4: Instance to concept alignment as seen in the embeddings of each model. The instances are colored according to the concept with the highest cosine similarity (after normalizing the similarities per concept with the $L_{2}$ norm). The most similar concept to an instance can be interpreted as the concept that is most closely aligned to the instance’s socioeconomic outcome.
@@
Not imposing any constraints on the latent space during model training is challenging for post-hoc concept interpretability as the resulting model might yield ambiguous CAVs that are hard to interpret [8]. As shown in the previous section, the Rank-N-Contrast pretraining procedure ensures that the embeddings lie on a manifold continuously ordered according to the target value. This impacts the learned CAVs as the concept examples are also projected along this manifold. Further, the high linear separability of the concepts on this manifold shown in Figure 3 suggests that the CAVs point to distinctive ranges of socioeconomic outcomes. We verify this by performing a similar analysis to [17] that identifies the most similar concept to each instance in the socioeconomic datasets based on the cosine similarity of the instance embeddings and the CAVs. The graphs on the left-hand side of Figure 4 reveal that in the contrastive pre-trained encoder, the concepts are clustered on typically non-overlapping intervals of the target values. In detail, for household income, the lower income areas are associated with the natural concepts of vegetation and water. Next, middle and high-income areas are aligned with sparse residential areas and agriculture fields while those with the highest income resemble medium residential areas. Regarding liveability, dense residential areas characterize the highest portion of the liveability range, particularly for the low and medium-liveable areas. Subsequently, the highest liveable areas are closely aligned with the natural concepts of water and vegetation. In summary, these results show that the Rank-N-contrast pretraining procedure improves the interpretability of the concept explanations as it enables one to intuitively associate the concepts to a continuous interval of values for the socioeconomic indicators. Conversely, the plots on the right-hand side of Figure 4 visualize that for both datasets, the models trained with the $L_{1}$ loss offer a limited understanding of how the concepts relate to the socioeconomic indicators as they typically do not form clear clusters in the latent space.
Insigths into Urban Planning with TCAV
Measuring the sensitivity of the concept to the model prediction with Equation (3) of Section 3.2.2 enables us to understand how the prediction changes when the visual patterns of the instances are perturbed in the direction of the concept. Furthermore, the concept clusters resulting from Rank-N-Contrast pretraining can unveil the correlation for intervals of socioeconomic indicators with the different urban and natural concepts. To shed light on the potential of such analysis for urban studies, in Figure 5 we visualize the magnitude of the TCAV values for the concept of vegetation. The left plot shows that perturbing the areas with low income in the direction of vegetation increases the model prediction for income. This positive effect decreases for areas with higher income. In other words, low-income areas can be seen rising in income as vegetation increases, with diminishing returns. The lowest-income areas in Paris can be found in the north and the northeast of the city, which is less green than most higher-income areas. For liveability, the opposite pattern is observed, where adding natural vegetation and water is only beneficial to the most liveable areas. While this may seem counterintuitive at first, it is important to consider that the liveability score is a composite socioeconomic outcome, comprising multiple dimensions at once. Neighborhoods in more densely populated areas often have better amenities and higher income opportunities. However, what is often lacking in these areas is access to nature. Conversely, while rural non-urban areas are abundant in natural vegetation, they are known to have lower incomes and access to amenities. The sensitivity plots of other scores can be found in Appendix, Section 9.
@@
Figure 5: The TCAV sensitivity of the vegetation concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization. For income, the sensitivity to vegetation is highest among the lowest income areas and it diminishes for the areas with higher income. In other words, adding vegetation to low-income areas increases the perceived income of the neighborhood by the model. For liveability, we observe the opposite effect, as the strongest increase in perceived liveability by the model in highly liveable areas can be achieved by increasing their amount of natural vegetation.
@@ 6 Conclusion
In this paper, we presented a pipeline that enables intuitive concept explanations of the deep learning models for socioeconomic outcome prediction. It consists of task-specific contrastive pretraining with the Rank-N-Contrast method and the TCAV approach for post-hoc concept testing. Our results show that the use of Rank-N-Contrast projects latent space embeddings in a continuously ordered manifold based on socioeconomic outcomes. This resulted in improved model performance for income prediction in Paris and yielded comparable results for liveability predictions in the Netherlands. The learned representation also contributed to intuitive concept explanations as we have shown that the concept activation vectors form clusters along continuous intervals of socioeconomic outcomes. For example, our analysis revealed that dense residential areas are typically associated with low liveability while more liveable areas are those that are close to the natural concepts of water and vegetation. The ordering of the concepts in the latent space is also beneficial for understanding their sensitivity to the model predictions. We illustrated this in Figure 5 which indicates that perturbing the instances of low income towards the vegetation concept improves their income score whereas for liveability this effect was observed for the highly liveable areas. Such insights open up the possibilities of our method for new insights into urban studies. Finally, we showed that our approach can be applied in different geographical locations as unlike concept bottleneck models, it does not require a set of concept labels specific to the location of the target study available during training.
@@ Acknowledgments
This work is jointly supported by the Munich Center for Machine Learning and by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab “AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond” (grant number: 01DD20001). We thank Dino Ienco, Nassim Ait Ali Braham, and Andrés Camero for the fruitfull discussions.
@@ 8 Concept Definitions
Table 2: Concept dataset composition from FLAIR Classes
@@ 9 Conceptual Sensitivities
Figures 6, 7, 8, 9, 10 and 11 visualize the conceptual sensitivities of the Rank-N-Contrast pre-trained encoder in both datasets for the concepts of water, agriculture, impervious, sparse residential, medium residential and dense residential, respectively.
Figure 6: The TCAV sensitivity of the water concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.
Figure 7: The TCAV sensitivity of the agriculture concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.
Figure 8: The TCAV sensitivity of the impervious surface concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.
Figure 9: The TCAV sensitivity of the sparse residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.
Figure 10: The TCAV sensitivity of the medium residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.
Figure 11: The TCAV sensitivity of the dense residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] with min-max normalization.

### v2

@@ Abstract
Predicting socioeconomic indicators from satellite imagery with deep learning has become an increasingly popular research direction. Post-hoc concept-based explanations can be an important step towards broader adoption of these models in policy-making as they enable the interpretation of socioeconomic outcomes based on visual concepts that are intuitive to humans. In this paper, we study the interplay between representation learning using an additional task-specific contrastive loss and post-hoc concept explainability for socioeconomic studies. Our results on two different geographical locations and tasks indicate that the task-specific pretraining imposes a continuous ordering of the latent space embeddings according to the socioeconomic outcomes. This improves the model’s interpretability as it enables the latent space of the model to associate concepts encoding typical urban and natural area patterns with continuous intervals of socioeconomic outcomes. Further, we illustrate how analyzing the model’s conceptual sensitivity for the intervals of socioeconomic outcomes can shed light on new insights for urban studies.
@@ 1 Introduction
In this work, we introduce a new expandability pipeline (visualized in Figure 1) that uses post-hoc concept explanations to reveal how deep learning models predict socioeconomic outcomes in terms of easy-to-understand concepts that represent typical urban and natural area patterns. It consists of three main steps: 1) a task-specific contrastive pre-training of a deep learning encoder to continuously order the latent space embeddings based on their socioeconomic outcome, 2) probing a linear layer on the top of the encoder to predict socioeconomic outcome, and 3) concept testing with the Testing with Concept Activation Vectors (TCAV) method [17]. We evaluate our pipeline on two tasks for estimating socioeconomic indicators from remote sensing imagery, namely income prediction in France [1], and liveability estimation in the Netherlands [21]. In summary, our work evaluates for the first time visual post-hoc concept-based explanations in remote sensing socioeconomic studies with the following main contributions:
@@ 2 Related Work
Although essential for their wider adoption in practice, the interpretability of the deep learning models for socioeconomic indicator estimation is rarely tackled [14]. The approaches in related studies mainly propose using saliency maps [1] or concept bottleneck models [25, 20, 21]. Concretely, Abitbol and Karsai train an EfficientNet deep learning model on aerial imagery to predict household income in five French cities [1]. They used the Grad-CAM saliency map method to infer the relevant regions for the model predictions. The identified regions are matched with the urban class polygons to reveal the relationship between urban classes and the income prediction of the deep learning model. Their results show that residential areas can both activate in case of high or low socioeconomic status depending on the city and area, whereas the presence of infrastructure such as motor- or railways is consistently related to a lower household income. When it comes to concept bottleneck models, they are used by Scepanovic et al. to predict the urban vitality in six Italian cities from Sentinel-2 imagery [25]. In the first step, the authors use a VGG-16 network and a convolutional autoencoder to predict six different land use and building block characteristics from image patches. These features are then used as input to the linear model that estimates urban vitality. The authors discover, that intersection density, small parks, anisotropicity, and land use mix positively influence vitality, while the variables block size and building height negatively impact the vitality. Further, Levering et al. train an end-to-end convolutional neural network with a semantic bottleneck to predict the landscape scenicness in the UK based on Sentinel-2 imagery [20]. The interpretable concepts in the semantic bottleneck layer encode the region’s land cover distribution. This architecture enables an explanation of the predicted scenicness score in terms of the land cover characteristics of the region. The authors find, that scenic areas are often associated with mountainous, coastal regions, while urban areas and the presence of human influence are negatively related to the landscape scenicness. The semantic bottleneck models are also used in [21] to predict the livability in the Dutch cities using aerial imagery. Again, the authors include a semantic bottleneck layer, which predicts five different intermediate domain scores. Their results indicate, that some intermediate domains such as buildings or physical environments are much better suited to be predicted from aerial imagery than others like safety or amenities.
@@ 4.3 Inference and Model Training
We predict the socioeconomic outcomes using a Resnet-50 encoder [15] for feature extraction followed by a linear layer that regresses the socioeconomic outcome. To train the model, we followed a similar training procedure as in [32]: First, we optimize the encoder for 400 iterations to minimize the RNC loss (the hyperparameter details are provided in Section 7 of the Appendix). Next, we freeze the weights of the feature encoder and train a single linear layer that uses the encoder output to predict the socioeconomic outcomes. The weights of the linear layer were optimized for 100 epochs with the $\mathcal{L}_{1}$ loss and we selected the model yielding the best $R^{2}$ score on the validation set. Our code is provided at github.com/IvicaObadic/rnc-4-visual-concept-explanations.
@@ 5.1 Prediction of Socioeconomic Indicators
To better understand the effect of task-specific pretraining on the latent space, in Figure 2 we use the t-distributed Stochastic Neighbor Embedding (t-SNE) method [28] to visualize the instance activations after the average pooling layer of the ResNet-50 encoder. The plots in the first column illustrate that the contrastive pre-trained encoder increasingly orders the instances in the latent space in an approximately 1D manifold according to their target label for both datasets. In contrast, the plots visualized in the second column depict that the models optimized to directly predict the regression score seem to result in a more scattered embedding space that fails to capture the continuous nature of the data, as there are multiple clusters of instances sharing similar target values located in various subspaces of the latent space. Finally, we note that for the task-specific pre-trained encoder, the liveability values typically display higher variance in their score compared to the income values for a local area in the embedding space. This can be attributed to the difference in complexity between these two tasks. The liveability reference data is based on 100 factors spanning 5 domains, from which liveability scores are derived [21]. As it is a composite score, the ranking difference between pairs will depend on more complex patterns, which might explain the comparable prediction results in Table 1 of the task-specific pretraining with the supervised trained encoder on the liveability task.
Figure 2: Household income instance activations in the average pooling layer visualized with t-SNE. By pre-training with Rank-n Contrast, the latent space can be ordered according to the task regression values, rather than the visual features of images. This results in an embedding space that aligns with the socioeconomic outcomes, and therefore is better suited for interpretability.
@@ 5.2 Concept Explanations
As described in Section 3.2, an essential step in TCAV is learning a concept activation vector (CAV) that encodes the direction in the latent space where the examples of a concept are located. In our study, we learn the CAVs per layer in the latent space of the ResNet-50 encoder by training a one-vs-all linear SVM classifier to discriminate the examples of the concept from randomly sampled 500 examples of all other concepts based on their layer activations. Figure 3 shows that while the concepts are already mostly linearly separable in the average pooling layer of the baseline encoder trained with the $L_{1}$ loss, enforcing a continuously ordered latent space with the Rank-N-Contrast pre-trained encoder further improves the concept discrimination. Notably, it enables consistently better linear separation for all concepts in the income task while for liveability the strongest improvements are observed in discriminating the sparse residential, medium residential, and dense residential concepts.
@@
Figure 4: Instance to concept alignment as seen in the average pooling layer of each model. The instances are colored according to the concept with the highest cosine similarity (after normalizing the similarities per concept with the $L_{2}$ norm). The most similar concept to an instance can be interpreted as the concept that is most closely aligned to the instance’s socioeconomic outcome.
@@
Not imposing any constraints on the latent space during model training is challenging for post-hoc concept interpretability as the resulting model might yield ambiguous CAVs that are hard to interpret [8]. As shown in the previous section, the Rank-N-Contrast pretraining procedure ensures that the embeddings lie on a manifold continuously ordered according to the target value. This impacts the learned CAVs as the concept examples are also projected along this manifold. Further, the high linear separability of the concepts on this manifold shown in Figure 3 suggests that the CAVs point to distinctive ranges of socioeconomic outcomes. We verify this by performing a similar analysis to [17] that identifies the most similar concept to each instance in the socioeconomic datasets based on the cosine similarity of the instance embeddings and the CAVs. The graphs on the left-hand side of Figure 4 reveal that in the contrastive pre-trained encoder, the concepts are clustered on typically continuous intervals of the target values. In detail, the left plot in the first row shows that lower to medium-income areas are associated with medium residential areas and impervious surfaces. Subsequently, the areas with higher-income typically relate to vegetation. Similarly, medium residential areas and vegetation also characterize the lower and the upper intervals of liveability, respectively. In contrast to the income task, the impervious surfaces occupy a larger portion of the liveability range, usually describing areas between the first and the third quartiles of liveability. In summary, these results show that the Rank-N-contrast pretraining procedure improves the interpretability of the concept explanations as it enables one to intuitively associate the concepts to a continuous interval of values for the socioeconomic indicators. Conversely, the plots on the right-hand side of Figure 4 visualize that for both datasets, the models trained with the $L_{1}$ loss offer a limited understanding of how the concepts relate to the socioeconomic indicators as they typically do not form coherent clusters in the latent space.
Insights into Urban Planning with TCAV
Measuring the sensitivity of the concept to the model predictions with Equation (3) of Section 3.2.2 enables us to understand how the prediction changes when the visual patterns of the instances are perturbed in the direction of the concept. This unveils whether the concepts contribute positively or negatively to the perceived socioeconomic outcome by the model. Furthermore, the concept clusters resulting from Rank-N-Contrast pretraining explain these correlations of the various urban and natural area concepts with intervals of the socioeconomic outcomes. This is exemplified in Figure 5 which shows the sensitivities of the instances to the vegetation concept. The sensitivities show the rural-urban gradient in income and liveability as the areas on the lower ends of both socioeconomic outcomes do not benefit much from adding vegetation, but its contribution and importance increase as income and liveability rise. Neighborhoods in more urbanized areas often have better amenities and higher wage expectations. However, access to nature is often lacking in these areas, which explains the positive contribution on the most positive ends of both socioeconomic outcome distributions. The sensitivity plots of other concepts shown in Section 9 of the Appendix illustrate the diverging trends that emerge for the concepts among socioeconomic outcomes. While income shows mostly clear linear trends, liveability displays a more complicated, non-linear relationship with the concepts. This may be explained by the fact that the liveability score is a complex composite score spanning many dimensions, whereas income reflects a single socio-economic outcome.
@@
Figure 5: The TCAV sensitivity of the vegetation concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and to the positive TCAV values, respectively. For income, the sensitivity to vegetation is highest among the high-income areas. In other words, adding vegetation to higher-income areas increases the perceived income of the neighborhood by the model. For liveability, we observe a similar effect, as the strongest increase in perceived liveability by the model in highly liveable areas can be achieved by increasing their amount of natural vegetation.
@@ 6 Conclusion
In this paper, we presented a pipeline that enables intuitive concept explanations of the deep learning models for socioeconomic outcome prediction. It consists of task-specific contrastive pretraining with the Rank-N-Contrast method and the TCAV approach for post-hoc concept testing. Our results show that the use of Rank-N-Contrast projects latent space embeddings in a continuously ordered manifold based on socioeconomic outcomes. This resulted in improved model performance for income prediction in Paris and yielded comparable results for liveability predictions in the Netherlands. The learned representation also contributed to intuitive concept explanations as we have shown that the concept activation vectors form clusters along continuous intervals of socioeconomic outcomes. For example, our analysis revealed that medium residential areas and impervious surfaces are associated with low to medium values of income and liveability. In contrast, the areas that lie in the upper socioeconomic outcome intervals display a strong association with the vegetation concept. Further, ordering the concepts in the latent space is also beneficial for understanding their sensitivity to the intervals of socioeconomic outcomes. We used this property to demonstrate the rural-urban gradient for the vegetation concept and the diverging trends of the other concepts among the socioeconomic outcomes. Such findings open up the possibility of using our method to gain new insights into urban studies. Finally, we showed that our approach can be applied in different geographical locations as unlike concept bottleneck models, it does not require a set of concept labels specific to the location of the target study available during training.
@@ Acknowledgments
This work is jointly supported by the Munich Center for Machine Learning and by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab “AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond” (grant number: 01DD20001). We thank Dino Ienco, Nassim Ait Ali Braham, and Andrés Camero for the fruitful discussions.
@@ 8 Concept Definitions
Table 2: Concept dataset composition based on the land cover segmentation maps provided in the FLAIR dataset [11].
@@ 9 Conceptual Sensitivities
Figures 6, 7, 8, 9, 10 and 11 visualize the conceptual sensitivities of the Rank-N-Contrast pre-trained encoder in both datasets for the concepts of water, agriculture, impervious surfaces, sparse residential, medium residential and dense residential, respectively.
Figure 6: The TCAV sensitivity of the water concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.
Figure 7: The TCAV sensitivity of the agriculture concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.
Figure 8: The TCAV sensitivity of the impervious surface concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.
Figure 9: The TCAV sensitivity of the sparse residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.
Figure 10: The TCAV sensitivity of the medium residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.
Figure 11: The TCAV sensitivity of the dense residential concept for the income (left) and liveability (right) datasets. The magnitude values are normalized in the range [-1, 1] by applying separate min-max normalization to the negative and the positive TCAV values, respectively.